# biblio.bib

### v1

@@ 2.4 Self-functionalization in vitro
We tested the efficiency of emulsion inversion in an in vitro set-up. We use agarose to mimic the mechanical properties of soft tissues, and inject droplets individually inside the gel in order to track the destabilization of the emulsion through fluorescence imaging. The injected droplets display diameters from 20 to , which is compatible with a range of in vivo applications.
We label the surface of the inner water droplets by grafting fluorescent streptavidin onto the biotinylated lipids, allowing us to measure the fluorescence on the outer surface of the oil droplet (see Fig 6). The functionalization is calculated by measuring the fluorescent signal intensity of a thick inner ring around the droplet, two hours after the injection (see Methods).
As expected, a higher concentration of lipids induces a higher emulsion stability and thus lowers functionalization of the oil droplet 2 hours after injection in medium. The functionalization obtained with a mix of of lipids and of Tween 20 is comparable to the functionalization obtained with only of lipids, but the addition of Tween lowers the surface enough to have droplets that are deformable enough in biological tissues (see Fig.5).
Considering the competing requirements of low surface tension ( ) and emulsion instability, the optimal formulation for achieving both deformability and self-functionalization is therefore identified as lipid, Tween 20.
@@
Next, we tested the drug delivery functionality of the inverted emulsion droplets. To do so, we checked if a local release of a loaded agent could be measured with this technique. Indeed, a very localized release of drugs or biological agents could be useful to decipher specific mechanisms during development by acting on the first neighboring cells in a specific region. To observe this effect, we loaded the aqueous phase of the inner droplets with large Dextran molecules () labelled with FITC. In this case, larger volumes of the aqueous phase can be used and the inverted emulsions can be prepared with PVDF filters (see Methods). The slower diffusion of this polymeric fluorophore (as compared with FITC alone for instance) allowed us to visualize some release events through which an inner droplet fuses with the outer surface, thus releasing a halo of fluorescence in the surrounding agarose gel (see Fig. 7).
@@ 2.5 Injection inside biological material
As a proof of concept, we tested the droplets in two different biological systems. First, we injected inverted emulsion droplets labeled with fluorescent streptavidin into dorsal forebrain organoids (see Methods). The organoids, that were embedded in agarose, were first imaged immediately after injection, showing fluorescent spherical droplets distributed within the tissue (Fig. 8A-B). A piece of agarose containing the organoid was then cut out and placed into a circular capillary. This procedure allowed to deform the organoid, which was further imaged through light sheet microscopy (see Methods). Elliptical droplets were clearly visible at approximately deep within the tissue (Fig. 8C-D), demonstrating that the droplets have sufficiently low interfacial tension for size droplets to be deform under mechanical stress within soft tissues such as organoids [ryuTransparentCompliant3D2021]. The local stress applied by the deformed tissue on the droplet can be estimated from its surface deformation. In particular, it is given by the formula , with and the maximum and minimum curvatures on the ellipsoidal droplet, respectively [Mongera2018]. For the droplet shown in Fig.8C, we can thus estimate a stress within the tissue .
Second, injections inside zebrafish embryos were performed at an early developmental stage (14-somites). The targeted area is near the eye and olfactory placode (OP), which is close to the surface and allows easy imaging. As shown in Fig. 8E, the droplets are well inserted in the embryos and their signal allows for their surface tracking in situ.
@@ 3 Methods
Oil and aqueous phases composition – The continuous oil phase is made of soybean oil (Sigma Aldrich) containing phospholipids 16:0 Biotinyl PE (Avanti Polar Lipids) and Tween 20 (Merck) at various concentrations. Briefly, a stock solution of phospholipids dissolved in chloroform was transferred in a glass vial at the chosen concentration. The solvent was then evaporated under nitrogen before of soybean oil containing Tween 20 was added to the dried lipids. The mixture was sonicated for 30 min at room temperature. On the hand, when Tween 20 was added, it was directly dissolved in soybean oil at concentrations ranging of r .
@@
Water-in-oil (W/O) emulsion preparation – Two luer-lock syringes (Fisher scientific) were filled with (syringe A) and (syringe B) of freshly vortexed surfactant-containing oil. After removal of any air bubble, syringe A was connected to 5 m pore size hydrophobic SPG pumping membrane connector (SPG Technology) and the oil was pushed through the connector. A 18 G needle (Terumo) was mounted on syringe B. About of oil was coarsely emulsified in the 50 l of aqueous phase through back and forth movements (5 to 10 times) inside the needle of syringe B. This coarse emulsion was then aspirated into syringe B before it was mounted onto the other side of the SPG membrane connector, while still making sure that no air bubbles are trapped in the syringe. The water droplet of coarse emulsion was then pushed back and forth through the connector, going from one syringe to the other. A total of five passages through the SPG membrane were completed, leading to the inverted emulsion used in the next stages. After each emulsion preparation, the SPG membrane connector was cleaned by flushing it with of Mili-Q water, followed by of soybean oil containing 16:0 Biotinyl PE.
@@
Droplet size distribution – After preparing the emulsion with the protocols detailed above, a few are inserted between a glass slide and a coverslip. The emulsion sample is then imaged through spinning disk confocal microscopy with a 60X oil immersion objective (1.4 N.A.) (see typical image in 3A). The droplet size was measured by applying a manual threshold and the Analyze Particles function of ImageJ. The diameter of droplets was then calculated from the measured area, assuming they have a spherical shape (90% of droplets have a measured circularity above 0.75) and the surfacic cumulative distribution function of droplet diameters was plotted (Fig. 3B), discarding all diameters measured below the microscope resolution (taken as ).
@@
Microinjection in agarose – The microinjection of inverted emulsion droplets was performed under a binocular microscope (Olympus, SZX2-FOFH). Borosilicate standard wall without filament capillaries of 0.69 mm inner and 1.2 mm outer diameter (Harvard Apparatus, 30-0042) were pulled using a micropipette puller (P-97, Sutter Instrument) under optimized settings (pressure = 500, heat=600, pull=120, velocity = 50, time = 225). to produce long, fine and initially sealed tips. Approximately of freshly prepared inverted emulsion was loaded into the capillary using microloader pipette tips (Eppendorf). The capillary tip was then manually cut using scalpel, achieving openings of a few micrometers in diameter. The capillary was finally mounted onto a holder connected to the pressure control system (Femto Jet 4i, Eppendorf).
@@
A solution of 0.5% (w/v) low-melting agarose (Sigma Aldrich) gel, prepared in 10 mM Tris buffer was poured in 35 mm glass bottom dish (Ibidi). The injection was performed at an intermediary stage, before full gelling was achieved. The microcapillary tip was positioned as close as possible to the bottom of the dish, in order to rethe main withing the working distance of the inverted microscope used for observation of self-functionalization dynamics. After injection, 1 mL of 10 mM Tris buffer were gently added on top of the gel to avoid evaporation effects and the dish was additionally sealed with its lid and wrapped with parafilm.
@@
Microinjection in zebrafish embryos – Using the same capillaries and injection protocol, droplets of inverted emulsion were injected in zebrafish embryos mounted in 0.5% low melting agarse at 14-somite stage (corresponding to 16 hpf). The embryos are obtained from a transgenic zebrafish line TgBAC(lamC1:lamC1-sfGFP) expressing Laminin gamma1 fused to sfGFP, under the control of its own regulatory sequences, as described in [Yamaguchi2022].
@@
Microinjection in organoids – Droplets were injected into a dorsal forebrain organoid derived from human induced pluripotent stem cells (hiPSCs). The organoid was generated and maintained in long-term culture following the protocol described in [Sloan2018], with adaptations from [Yoon2019]. For quality control, routine mycoplasma testing, SNP analysis, and immunohistochemistry were performed to ensure genomic integrity and correct brain region specification. The hiPSC line used was the UKBi013-A-GCaMP6f iPSC reporter line, derived from the parental line UKBi013-A (https://hpscreg.eu/cell-line/UKBi013-A) via CRISPR/Cas9-mediated genome editing, as described in detail in [Li2025]. At the time of droplet injection, the organoid was 100 days old. Samples were embedded in 1.5% low-melting-point agarose (Invitrogen) prepared in sample medium and loaded into glass capillaries (Drummond, Wiretrol II). The capillary was then placed into the imaging chamber, which was filled with sample medium. Prior to imaging, the sample was gently extruded from the capillary to ensure direct exposure to the light-sheet.
Interfacial tension measurements – Interfacial tension between oil and water were measured by rising drop tensiometry. A blunt needle (G 23, PHYMEP) was inserted through the bottom of a polystyrene macro cuvette (Fisher scientific), filled with an aqueous solution of 10mM Tris (pH = 7.4 ). The oil phase (or inverted emulsion) was injected through the needle until a rising pear shaped droplet was formed (see Fig. 5A-C). The set-up was illuminated in transmission thanks to a led panel and the droplet was imaged with a camera over time (an image every 5 minutes during 20 to 40 minutes).
The interfacial tension was estimated from the axisymmetric shape profile of the rising drop using Pendant Drop plugin in Fiji [Daerr2016].
From each condition, we get the mean measurement of surface tension at each time point, which we fit with an exponential decay to a finite value function. This finite value obtained from the fit is our reported equilibrium value in Fig 5D-E .
Imaging – The inverse emulsion was imaged through spinning disk confocal microscopy (Spinning Disk Xlight V2, Gataca systems). Droplets in agarose as well as underformed organoids were imaged through epifluorescence microscopy (inverted Nikon Eclipse Ti2-E), using a Nikon, PLAN APO Lambda D, 20x/0.8 objective. In order to track the droplets over time, z-stacks were acquired to compensate for droplet drifts. Image analysis was then performed in the identified equatorial plane of each droplet.
Light-sheet fluorescence imaging was performed using a previously described setup [Hubert2023]. Two continuous-wave lasers were used for excitation: a 561 nm laser (Vortran) for imaging red fluorescence and a 488 nm laser (Oxxius) for imaging green fluorescence. The light-sheet was generated via a galvanometric mirror (Thorlabs, GVS001) and directed into the sample through an illumination objective (Zeiss, 5X A-Plan, NA=0.12). Fluorescence emission was collected using a detection objective (Leica, HC FLUOTAR L 25x/0,95W VISIR) and a tube lens (Olympus), resulting in a total system magnification of 22.5X. An appropriate emission filter was placed at the detection path depending on the fluorescence channel being recorded. Images were acquired using a Hamamatsu camera. The total illumination power delivered to the sample was kept below 1 mW.
@@
Image analysis – Water phase droplets in the emulsion (Fig.3) was measured by first applying an intensity threshold by hand that ensured all droplets that were visible were detected. Then, a watershed function is applied to separate the few droplets that are too close and could appear merged by the thresholding step. Last, the shape and size of all particles detected this way was measured, removing all particles that are smaller that 6 pixels or that have a circularity lower than 0.5 to remove potential noise. In practice, less than 1% of particles detected were in these criteria, which indicates that the droplets signal and detection algorithm was accurate.
Droplet functionalization analysis (Fig.6) was done in three major steps. First, the droplet images were prepared for contour detection with saturation and bleaching compensation. The bleaching compensation is applied first by running a histogram-matching algorithm on series of the same field of view over time. Then, the saturation is done by setting all pixels that have an intensity higher than 10 times the intensity of the background. From experience, this setup ensured that no pixel of interest (from the droplet contour) were saturated, but most pixels from the very bright inner droplet spots were saturated. Second, the droplet contour was detected by hand, with the user pointing at 3 pixels that are on the droplet edge. For droplets with a very bright contour (which means a very high functionalization), an automatic contour detection algorithm was used and it resulted in no difference with the contour detected by the user. Last, the normalized intensity of each droplet (which is the final proxy for functionalization) is calculated by measuring the mean pixel intensity in a ring which exterior is the outer edge of the droplet, and dividing this measured value by the mean background intensity.

### v3

@@ 2.4 Self-functionalization in vitro
We use agarose to mimic the mechanical properties of soft tissues, and inject oil droplets individually inside the gel in order to track the destabilization of the emulsion through fluorescence imaging (see Methods). The injected oil droplets display diameters from 20 to , which is compatible with a range of in vivo applications.
We label the surface of the inner water droplets by grafting fluorescent streptavidin onto the biotinylated lipids.The fluorescence is initially visible only in the center of the oil droplet, where the inner water droplets are located (see Fig.7A, left panel). Over time, a signal appears on the edge of the injected oil droplet (Fig.7A, middle and right panels), indicating that some of the inner water droplets have fused with the outer surface over time, allowing the fluorescent streptavidin to relocalize on this outer oil/water interface.
We quantify this functionalization process by measuring the fluorescent signal intensity of a thick inner ring around the oil droplet, two hours after the injection (see Methods). On the one hand, we find that a higher concentration of lipids slightly lowers the functionalization signal (see dark to light blue columns in Fig.7B). This is due to the fact that a higher lipid concentration tends to stabilize more the inner water droplets. On the other hand, the addition of Tween 20 at also lowers the functionalization signal (see pink column in Fig.7B), making it comparable to the highest lipid concentration (light blue) but with a lower surface tension. We do not explore the functionalization efficiency at higher Tween 20 concentrations, which would decrease its efficiency while only having a marginal effect on surface tension (Fig.5B).
Considering the competing requirements of low surface tension ( ) and functionalization, we choose lipid, Tween 20 as a formulation that balances well the given constraints.
@@
Next, we tested the drug delivery functionality of the oil droplets. To do so, we checked if a local release of a loaded agent could be measured with this technique by loading the inner water droplets with large Dextran molecules () labelled with FITC. In this case, because the aqueous phase is made of less expensive and rather robust material, it can be prepared in larger volumes. Therefore the inverted emulsions is made with PVDF filters (see Methods). The slow diffusion of this polymeric fluorophore (as compared with FITC alone for instance) allowed us to visualize some release events through which an inner water droplet fuses with the outer surface, thus releasing a halo of fluorescence that diffuses in the surrounding agarose gel (see Fig.8, left to right panels).
In all these tests of functionality, the chosen molecules and fluorescent labels are not surface active agents. Therefore they do not affect the surface tension of the inner water droplets. However, it is important to bear in mind that the encapsulation of surfactants could decrease the self-functionalizing properties of the oil droplet. For instance, the encapsulation of fluorescent colloids did strongly stabilize the inner water droplets oil/water interface, akin to pickering emulsions, and hindered the self-functionalizing properties of the sensors (data not shown).
2.5 Mechanical characterization of the sensors
Oil droplets have been used inside biological tissues as stress probes [29]. In this case, if the oil droplets are embedded in the tissues, the accessible measurement is their local surface deformations. In turn, using the Laplace balance on the oil droplet’s surface lead to a measurement of the anisotropic stresses at every point of its contour [30]. Alternatively, these deformations can be approximated in 2D by an elliptical shape, allowing to measure more simply the anisotropic stress from the difference between its principal stresses. In this case, the stress anisotropy on the oil droplet is given by the formula , with and the maximum and minimum curvatures on the ellipsoid, b and a being its long and short semi-axis respectively. While the Laplace balance is a mechanical equilibrium that holds regardless of constitutive laws of the surrounding medium, if one doesn’t know the modulus of the surrounding tissue it is not possible to relate an externally applied force on the tissue to the observed oil droplet deformation. In other words it is used as a local rather than global probe. However, it can still be a powerful proxy for local stress anisotropies in tissues, as shown in [30].
On the other hand, knowing the elastic modulus of the surrounding medium, one can compute the expected deformation of the oil droplet as a function of an externally applied stress, extending the Eshelby framework to the situation of a liquid inclusion with a surface tension [39]. Therefore, we probed our sensors in a model mechanical excitation set-up [37, 40] (see Fig. 9A). Briefly, oil droplets are injected inside agarose gels mounted inside spectroscopy cuvettes. A piston is positioned on top of the gel so that it can apply a known force while the deformation of the oil droplets are visualized by a camera positioned laterally to the cuvette. At each force step, their deformation is measured by fitting an ellipse to their contour, as shown in Fig. 9B.
This approach allowed us to measure the oil droplet aspect ratio and compare it with the theoretical model proposed by Style et al. [39], assuming a Poisson’s ratio of , according to the following relation:
where is the uniaxial stress applied to the gel, is the Young’s modulus of the gel, and is the interfacial tension between the oil droplet and the gel. Equation 1 shows that and can be different, depending on , and , even if is the expected stress anisotropy in the gel in the absence of droplet. In order to probe different regimes, experiments were performed using two agarose gels: 0.6% agarose with a measured Young’s modulus of , and 1% agarose with . We injected oil droplets prepared either as inverted emulsions containing lipid and Tween 20, or as pure oil droplets without surfactant. The latter, associated with a higher surface tension, were used in softer gels in order to probe a regime in which . In that case, the expected surface tension of the oil droplet is [41] [42].
Figure 9C shows the measured aspect ratio of the oil droplets as a function of . As predicted by equation 1, large and softer oil droplets embedded in stiff gels for which (squares in Fig. 9C) exhibit an aspect ratio independent of their size (note the large spread of sizes, ranging from –). Conversely, when (circles in Fig. 9A), surface tension resists deformation, resulting in a smaller aspect ratio compared to that of an infinitely deformable oil droplet.
From equation 1, the aspect ratio of the oil droplet is expected to depend only on the dimensionless parameter
that contains information on its size, surface tension and gel stiffness. As shown in Fig. 9D, all data collapse onto a single curve when plotted as a function of . This analysis confirms the importance of surrounding material properties, together with the size of the oil droplets and their surface tension, to understand their deformation inside a material subject to mechanical load. In other words, if the material properties of the surrounding tissue and sensor are known, the global stress applied on the tissue can be derived through the above-described framework. If only the sensor properties are known, its use is limited to probing local anisotropic stresses on the oil droplet.
@@ 2.6 Injection inside biological material
As a proof of concept, we tested the oil droplets in two different biological systems. First, we injected inverted emulsion oil droplets labeled with fluorescent streptavidin into dorsal forebrain organoids (see Methods). The organoids, that were embedded in agarose, were first imaged immediately after injection, showing fluorescent spherical oil droplets distributed within the tissue (Fig. 10A-B). A piece of agarose containing the organoid was then cut out and placed into a circular capillary. This procedure allowed to deform the organoid, which was further imaged through light sheet microscopy (see Methods). Elliptical oil droplets were clearly visible at approximately deep within the tissue (Fig. 10C-D), demonstrating that they have sufficiently low interfacial tension for oil droplets to be deform under mechanical stress within soft tissues such as organoids [43]. Since in this case the absolute modulus of the tissue is unknown, we compute the local anisotropic stress on the oil droplet through the formula , as described in the previous section. For the oil droplet shown in Fig.10C, we can thus estimate an anisotropic stress on the droplet (uncertainty calculated from the image resolution). This value compares well with previously reported values measured with solid sensors in deformed cellular aggregates [28].
Second, injections inside zebrafish embryos were performed at an early developmental stage (14-somites). The targeted area is near the eye and olfactory placode (OP), which is close to the surface and allows easy imaging. As shown in Fig. 10E, the oil droplets are well inserted in the embryos and their signal allows for their surface tracking in situ. Since this area of the embryo is mostly composed of extracellular matrix and loosely-arranged migrating neural crest cells as well as other cell types of the peri-occular mesenchyme [44], the oil droplet does not exhibit any measurable deformation. Conversely, when the sensors are injected in between two compact tissues such as the retina and the lens, as in Fig. 10F, their deformation is significant and measurable. Furthermore, the deformation of the oil droplet can be tracked over time, leading to an estimation of the local anisotropic stress evolution on the droplet during development in this particular area of the zebrafish embryo (see Fig.10G). The decrease of local anisotropic stress could be assigned to either relaxation of local tissue stresses through viscous dissipation or to an increase of the stress orthogonal to the long axis of the droplet.
@@ 3 Methods
Oil and aqueous phases composition – The continuous oil phase consisted of soybean oil (Sigma-Aldrich) containing 16:0 Biotinyl PE phospholipids (Avanti Polar Lipids) and Tween 20 (Merck) at various concentrations. Briefly, a stock solution of phospholipids dissolved in chloroform was transferred to a glass vial at the desired concentration. The solvent was then evaporated under a nitrogen stream, after which of soybean oil containing Tween 20 was added to the dried lipids. The mixture was sonicated for 30 minutes at room temperature. When Tween 20 was used, it was directly dissolved in soybean oil at final concentrations of or .
@@
Water-in-oil (W/O) emulsion preparation – Two luer-lock syringes (Fisher scientific) were filled with (syringe A) and (syringe B) of freshly vortexed surfactant-containing oil. After removal of any air bubble, syringe A was connected to pore size hydrophobic SPG pumping membrane connector (SPG Technology) and the oil was pushed through the connector. A 18 G needle (Terumo) was mounted on syringe B. About of oil was coarsely emulsified in the of aqueous phase through back and forth movements (5 to 10 times) inside the needle of syringe B. This coarse emulsion was then aspirated into syringe B before it was mounted onto the other side of the SPG membrane connector, while still making sure that no air bubbles are trapped in the syringe. The water droplet of coarse emulsion was then pushed back and forth through the connector, going from one syringe to the other. A total of five passages through the SPG membrane were completed, leading to the inverted emulsion used in the next stages. After each emulsion preparation, the SPG membrane connector was cleaned by flushing it with of Mili-Q water, followed by of soybean oil containing 16:0 Biotinyl PE.
@@
In both cases (SPG membrane or filter), this emulsification step can be completed in a few minutes. The prepared emulsion should then be used immediately for injection as detailed below.
@@
Inner water droplets imaging – After preparing the emulsion with the protocols detailed above, a few are inserted between a glass slide and a coverslip. The emulsion sample is then imaged through spinning disk confocal microscopy with a 60X oil immersion objective (1.4 N.A.) (see typical image in Fig. 3A).
Interfacial tension measurements – Interfacial tension between oil and water were measured by rising drop tensiometry. A blunt needle (G 23, PHYMEP) was inserted through the bottom of a polystyrene macro cuvette (Fisher scientific), filled with an aqueous solution of 10mM Tris (pH = 7.4 ). The oil phase (or inverted emulsion) was injected through the needle until a rising pear shaped oil droplet was formed (see 5A-D). The set-up was illuminated in transmission thanks to a led panel and the oil droplet was imaged with a camera over time (an image every 5 minutes during 20 to 40 minutes).
The interfacial tension was estimated from the axisymmetric shape profile of the rising drop using Pendant Drop plugin in Fiji (see 5B) [45].
From each condition, we get the mean measurement of surface tension at each time point, which we fit with an exponential decay to a finite value function (see 5E). This finite value obtained from the fit is our reported equilibrium value in Fig 6A-B .
@@
Microinjection in agarose – The microinjection of inverted emulsion oil droplets was performed under a binocular microscope (Olympus, SZX2-FOFH). Borosilicate standard wall without filament capillaries of 0.69 mm inner and 1.2 mm outer diameter (Harvard Apparatus, 30-0042) were pulled using a micropipette puller (P-97, Sutter Instrument) under optimized settings (pressure = 500, heat=600, pull=120, velocity = 50, time = 225). To produce long, fine and initially sealed tips. Approximately of freshly prepared inverted emulsion was loaded into the capillary using microloader pipette tips (Eppendorf). The capillary tip was then manually cut using scalpel, achieving openings of a few micrometers in diameter. The capillary was finally mounted onto a holder connected to the pressure control system (Femto Jet 4i, Eppendorf). To ensure constant inverted emulsion volume injection, two approaches can be used: either the glass capillaries should be pre-opened with the reproducible tip opening, followed by standardized microinjector parameters; or if the capillary tips are less controlled in size, the microinjector can be calibrated in advance for each capillary, in order to deliver the desired volume. Note that cheaper manual injectors can also be used (such as Eppendorf celltram oil pumps) but require a lot more practice by the users in order to reach precise volumes.
@@
A solution of 0.5% (w/v) low-melting agarose (Sigma Aldrich) gel, prepared in 10 mM Tris buffer was poured in 35 mm glass bottom dish (Ibidi). The injection was performed at an intermediary stage, before full gelling was achieved. The microcapillary tip was positioned as close as possible to the bottom of the dish, in order to remain within the working distance of the inverted microscope used for observation of self-functionalization dynamics. After injection, 1 mL of 10 mM Tris buffer were gently added on top of the gel to avoid evaporation effects and the dish was additionally sealed with its lid and wrapped with parafilm.
Mechanical characterization of stress sensor – To probe the stress-sensing properties of the oil droplets, we used a mechanical excitation set-up, described in detail in ref. [37, 40], with slight modifications. Briefly, a gel slab containing the oil droplets can be lifted vertically using a manual micrometric Z stage to indent a fixed piston, itself mounted on a force measurement setup. The force sensor consists in a dual cantilever beam (whose stiffness has been independently measured). The deflection of the cantilever is measured with a capacitive sensor, from which the normal is deduced (range ), with a measurement noise of about . In order to work with a cube of agarose gel the bottom of the cuvette was filled with polydimethylsiloxane (PDMS). Moreover, two thin spacers made of Plexiglas (thickness ) can be placed on two opposing sides of the cuvette, enabiling, after their removal, lateral expansion of the gel during compression. Agarose gels at concentrations 1% and 0,6% were used to tune the corresponding elastocapillary length. Freshly prepared inverted emulsion was injected in 1% agarose gel, as described above, before full gelling was achieved. The applied force was increased in steps, starting from 0 N up to maximum of , while monitoring the shape of oil droplets with a Pointgrey camera (BFLY-U3-23S6C-C) facing a side of the cuvette. Once the maximal force was reached, the cuvette was shifted downward to reduce the load to , allowing oil droplets to relax and go back to their spherical shape. For the 0.6% agarose gel, soybean oil droplets were injected instead, and the applied force was increased in steps from up to , following the same protocol.
@@
Microinjection in zebrafish embryos – Using the same capillaries and injection protocol, oil droplets of inverted emulsion were injected in zebrafish embryos mounted in 0.5% low melting agarse at 14-somite stage (corresponding to 16 hpf). It should be noted that depending on the tissue accessibility and mechanical properties, glass capillaries used for injection can be tailored by adjusting the micropipette puller settings. For example, if the target site is located superficially underneath the stiff structure, such as skin, shorter capillaries are preferable as they are less prone to bending and provide better penetration. In contrast, if the target is located deeper, within softer tissue, usually longer and thinner capillaries are required, in order to reach the site effectively and not cause damage. The embryos are obtained from a transgenic zebrafish line TgBAC(lamC1:lamC1-sfGFP) expressing Laminin gamma1 fused to sfGFP, under the control of its own regulatory sequences, as described in [46].
@@
Microinjection in organoids – Oil droplets were injected into a dorsal forebrain organoid derived from human induced pluripotent stem cells (hiPSCs). The organoid was generated and maintained in long-term culture following the protocol described in [47], with adaptations from [48]. For quality control, routine mycoplasma testing, SNP analysis, and immunohistochemistry were performed to ensure genomic integrity and correct brain region specification. The hiPSC line used was the UKBi013-A-GCaMP6f iPSC reporter line, derived from the parental line UKBi013-A (https://hpscreg.eu/cell-line/UKBi013-A) via CRISPR/Cas9-mediated genome editing, as described in detail in [49]. At the time of injection, the organoid was 100 days old. Samples were embedded in 1.5% low-melting-point agarose (Invitrogen) prepared in sample medium and loaded into glass capillaries (Drummond, Wiretrol II). The capillary was then placed into the imaging chamber, which was filled with sample medium. Prior to imaging, the sample was gently extruded from the capillary to ensure direct exposure to the light-sheet.
Imaging – The oil in water emulsion was imaged through spinning disk confocal microscopy (Spinning Disk Xlight V2, Gataca systems) using a 60X oil immersion objective. Oil droplets in agarose and underformed organoids were imaged through epifluorescence microscopy (inverted Nikon Eclipse Ti2-E), using a Nikon, PLAN APO Lambda D, 20x/0.8 objective. In order to track the oil droplets over time, z-stacks were acquired to compensate for drifts. Image analysis was then performed in the identified equatorial plane of each oil droplet.
Light-sheet fluorescence imaging was performed using a previously described setup [50]. Two continuous-wave lasers were used for excitation: a 561 nm laser (Vortran) for imaging red fluorescence and a 488 nm laser (Oxxius) for imaging green fluorescence. The light-sheet was generated via a galvanometric mirror (Thorlabs, GVS001) and directed into the sample through an illumination objective (Zeiss, 5X A-Plan, NA=0.12). Fluorescence emission was collected using a detection objective (Leica, HC FLUOTAR L 25x/0,95W VISIR) and a tube lens (Olympus), resulting in a total system magnification of 22.5X. An appropriate emission filter was placed at the detection path depending on the fluorescence channel being recorded. Images were acquired using a Hamamatsu camera. The total illumination power delivered to the sample was kept below 1 mW.
@@
Image analysis – The size distribution of the inner water droplets (as shown in Fig.3) was measured on confocal images acquired with a 60X oil immersion objective and with parameters that prevented any saturation. First, a manual intensity threshold was applied to ensure the detection of all visible water droplets. Then, a watershed function was performed to separate the few water droplets that were too clustered and could appear merged by the thresholding step. Last, the shape and size of all particles detected this way was measured, removing all particles that are smaller than 6 square pixels, corresponding to a diameter of , or that have a circularity lower than 0.5 to remove potential noise. In practice, less than 1% of particles detected were in these criteria, which indicates that the signal and detection algorithm was accurate. All diameters measured below the microscope resolution (taken as ) were assigned the value d=. These water droplets accounted for less than 5% of the detected ones and thus have a negligible effect on the mean size and standard deviation.
Oil droplet functionalization analysis (Fig.7) was done in three major steps. First, the images were prepared for contour detection with saturation and bleaching compensation. The bleaching compensation is applied first by running a histogram-matching algorithm on series of the same field of view over time. Then, the saturation is done by setting all pixels that have an intensity higher than 10 times the intensity of the background. From experience, this setup ensured that no pixel of interest (from the oil droplet contour) were saturated, but most pixels coming from the very bright inner water droplets were saturated. Second, the oil droplet contour was detected by hand, pointing at 3 pixels on its edge. For very bright contours (which means a very high functionalization), an automatic detection algorithm was used and it resulted in no difference with the manually detected contour. Last, the normalized intensity of each oil droplet (which is the final proxy for functionalization) is calculated by measuring the mean pixel intensity in a ring around its contour, and dividing this value by the mean background intensity.